Can you please solve this question for me: 50+(2*10)/4*2
Answer by gonzo(654)

order of solving is:
going from left to right
solve all expressions within parentheses first
-----
do exponentiation first
then mult/div
then add/subt
multiple occurrences of mult/div are done left one first.
multiple occurrences of add/subt are done left one first.
-----
your equation breaks down as follows:
-----
50 + (2*10) / 4 * 2 =
50 + 20 / 4 * 2 =
50 + 5 * 2 =
50 + 10 =
60
-----
the expression within the parentheses was done first.
since there was no exponentiation, the multiplication or division was done next going from left to right.
first was 2*10 within the parentheses which became 20
next was 20 / 4 which became 5
next was 5 * 2 which became 10
last was 50 + 10 which became 60
-----
answer would be 60.